OpenStudy (anonymous):

Help Simplify, rationalize denominators: 3 radical 18 + 2 radical 72

OpenStudy (anonymous):

waiting yyyyyyyyyyyaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaarrrrrrrrrrrrrrr

OpenStudy (anonymous):

\[3\sqrt{18} + 2\sqrt{72}\]

OpenStudy (saifoo.khan):

\[21\sqrt{2}\]

OpenStudy (anonymous):

how

OpenStudy (saifoo.khan):

Good question. lol.

OpenStudy (anonymous):

\[3\sqrt{18}+2\sqrt{72} = 3\sqrt{9}\sqrt{2}+2\sqrt{9}\sqrt{4}\sqrt{2} = 3x3\sqrt{2}+2x3x2\sqrt{2} = (9+12)\sqrt{2} = 21\sqrt{2}\]